OpenStudy (lilsis76):

MAJOR HELP!!!Trigonometric equations: Find all solutions of the equation in the interval 0° ≤ x ≤ 360°. a) sin 3 Theta = -squrt2 / 2 b) 2sin^2 Theta - cos 2Theta = 0 c) sin2Theta = squrt3 cos 2Theta...HINT: Divide by cos 2Theta

OpenStudy (jonnyvonny):

a) sin(x) is -sqrt(2)/2 at 240 and 300, so, you need to divide both of those by 3 to counter-act the multiplication; you'll yield same answer: sin (3(240/3)) = sin(45) 80, 100 b) Not too sure of. c) sin(2x)/cos(2x)=sqrt(3) Sin(2x)/cos(2x) is the same as tan(2x). tan(x)=sqrt (3) at pi/3 and 2pi/3= 60 and 120. Same principle as a; since we're multiplying; divide both by 2. 60/2=30 120/2=60 30, 60.

OpenStudy (jonnyvonny):

Sorry about reposting twice; I read it wrong and had to correct a typo.

OpenStudy (lilsis76):

so for a) i started to try and work it out and I see that \[\frac{ -\sqrt{2} }{ 2 } \] is 45 degrees. but its negative so the sine negative would be the 45 deg. in the quadrant 3 and 4. so it gave me 225 degree and 315 degree right? haha its okay, im trying to understand this whole thing

OpenStudy (lilsis76):

@JonnyVonny do i understand or did it right?? im going by what I thought haha sorry. Imma look at yours now

OpenStudy (lilsis76):

here ill draw it out maybe this would help. @JonnyVonny

OpenStudy (lilsis76):